Write down expression of work done in isothermal process.
1
Expert's answer
2020-10-12T07:51:09-0400

Solution:


W = integral of p dV.

If we are working with an ideal gas, then

p ="\dfrac{nRT}{V}" and

W = integral of nRT "\dfrac{dV}{V}"= nRT ln("\dfrac{V_{2}}{V_{1}}").

If we are not working with an ideal gas, we will have to supply more information.

Another possible answer is W = Q (because there is no internal energy change), but that's probably not what they're after.